OpenStudy (ohmybookness):

solve algebraically: x^2 +y^2 +2x-4y=20 3x+4y=5

OpenStudy (ohmybookness):

@Awolflover1 @BioHazard9064 @countrygirl1431 @dan815 @eliassaab @F_Jayyy @gottennis121 @HelloKitty17 @imqwerty @jabez177 @KendrickLamar2014 @Learner11 @mathstudent55 @nincompoop @ocormican @phi @Qwertty123 @trwatkins1 @unicwaan @Vitgan13

OpenStudy (ohmybookness):

@Wendy.Ivette11714 @xXsebastianXx @YoungStudier @victoria_grace @Zeronknight

OpenStudy (ohmybookness):

I know the answer I just need to algebraically solve it :/

OpenStudy (lynfran):

we complete the square on x^2+y^2+2x-4y=20 we first group the corresponding variable so x^2+2x+y^2-4y=20 what we do now is, take half of the 2 infront of the x and half of the -4 and then..square them... add the squared value to the right side of the equation (x+1)^2+(y-2)^2=1+4+20

OpenStudy (ohmybookness):

so from there what do i do?

OpenStudy (ohmybookness):

someone help please

OpenStudy (lynfran):

(x+1)^2+(y-2)^2=25 so x=-1 , y=2 with r=5 since its a circle with radius 5 and center (-1,2) are those ur answer

OpenStudy (ohmybookness):

here Ill show you the file. It would be #8

OpenStudy (lynfran):

o they are grouped equations.. taught they were separate

OpenStudy (ohmybookness):

yeah they're grouped. I thought the same at first

OpenStudy (lynfran):

ok i had to actually work this out on paper

OpenStudy (lynfran):

let me retake the picture clearer

OpenStudy (ohmybookness):

Wow!! thank you so much!

OpenStudy (ohmybookness):

y answer correct though?

OpenStudy (ohmybookness):

was my*

OpenStudy (lynfran):

yes the answers were correct ... actually those 2 points are the points of intersection where those 2 equations meet